f(x+h)−f(x)=7hx2−5hx+1h2x+4h2+7h3.
Find f′(x).
1
Expert's answer
2020-11-09T14:57:27-0500

"\displaystyle f(x+h)−f(x)
=7hx^2−5hx+h^2x+4h^2+7h^3 \\


\frac{f(x+h)−f(x)}{h}
=7x^2−5x+hx+4h +7h^2\\


\begin{aligned}
\lim_{h \rightarrow 0} \frac{f(x+h)−f(x)}{h} &= \lim_{h \rightarrow 0}(7x^2−5x+hx+4h +7h^2) 
\\&= 7x^2 - 5x + 0 + 0 + 0 = 7x^2 - 5x
\end{aligned}\\


\therefore f'(x) = 7x^2 - 5x"